OpenStudy (anonymous):

A block of mass m is at rest on a 20° slope. The coefficient of static friction between the box and the floor is 0.64. The block is connected by a massless string over a massless, frictionless pulley to a hanging block of mass 2.0 kg. What is the minimum mass m that will stick to the surface and not rise up the slope?

OpenStudy (zephyr141):

as always draw your Full Body Diagram out. then label the forces acting on each object.

OpenStudy (zephyr141):



OpenStudy (zephyr141):

for m2 do this\[y:T-m_2g=0\]\[y:T=m_2g\]then for the block on the incline\[y:n-m_1gsin20=0\]\[y:n=m1gsin20\]and for x\[x:T-f=0\]\[x:T=f\]\[x:T=\mu n\]

OpenStudy (zephyr141):

\[m_2*g=\mu *m_1*g*\sin20\]put in your known values and solve for m1. i think this is how you do it.

OpenStudy (zephyr141):

\[m_2=\mu m_1\sin20\]

OpenStudy (zephyr141):

g could be divided out.

OpenStudy (zephyr141):

damn i forgot one force. hold on.

OpenStudy (zephyr141):

ok it should be this...\[m_1=\frac{m_2}{\mu \sin20+\cos20}\]